One angle is twice its supplement increased by 15 degrees. Find the measure of the two supplementary angles.
Answer by nerdybill(7384)

One angle is twice its supplement increased by 15 degrees. Find the measure of the two supplementary angles.
.
Let x = supplement of angle
then
2x+15 = angle
.
x + 2x+15 = 180
3x + 15 = 180
3x = 165
x = 55 degrees
.
other angle:
2x+15 = 2(55)+15 = 110+15 = 125 degrees
.
answer: 55 and 125 degrees